OpenStudy (sleepyjess):

Verify the identity. \(\sf\dfrac{cos~x}{1+sin~x}~+~\dfrac{1+sin~x}{cos~x}~=~2~sec~x\)

ganeshie8 (ganeshie8):

you figured it out ? :)

OpenStudy (sleepyjess):

No

ganeshie8 (ganeshie8):

you may start by combining both the fractions on the left hand side

OpenStudy (sleepyjess):

I would have to have a common denominator so, \(\sf\dfrac{(cos~x)(1~+~sin~x)}{(1~+~sin~x)(cos~x)}~+~\dfrac{(1~+~sin~x)(cos~x)}{(cos~x)(1~+~sin~x)}\) \(\sf\dfrac{cos x + sin x cos x}{cos x + sin x cos x}~+~\dfrac{cos x + sin x cos x}{cos x + sin x cos x}\) \(\sf\dfrac{2(cos x + sin x cos x)}{cos x + sin x cos x}\)

OpenStudy (sleepyjess):

@e.mccormick

OpenStudy (e.mccormick):

Hmmm.... Your representations of 1 to do common denomintor seem wrong. \(\sf\dfrac{cos~x}{1+sin~x}~+~\dfrac{1+sin~x}{cos~x}~=~2~sec~x\) \(\sf\dfrac{cos~x}{1+sin~x}\cdot\dfrac{cos~x}{cos~x}~+~\dfrac{1+sin~x}{cos~x}\cdot\dfrac{1+sin~x}{1+sin~x}~=~2~sec~x\)

OpenStudy (e.mccormick):

They key is that anything divided by itself is 1 and therefore does not change the fraction. \(\dfrac{1}{4}=\dfrac{1}{4}\cdot\dfrac{a}{a}\) So to get a cos x onto the bottom of the left most fraction, I multiply by 1 in the form of \(\dfrac{\cos x}{\cos x}\). Similar rules for doing the other fraction to get the common denominator.

OpenStudy (e.mccormick):

Do you see how fixing that mistake will lead to a different answer?

OpenStudy (sleepyjess):

Yes

OpenStudy (e.mccormick):

One of the key things to remember in math is that there are two basic operations that do not change things. First is the adding of 0 and second is multiplying by 1. So why is this useful? Well, let me look at them each: +42-42 = 0, therefore you can add +42-42 to anything and it will not change the value of an equation. If you need to factor something with a +42, that could make it more workable. You use this basic tactic in completing the square. Multiplying by 1 I just showed. If it is 1, \(\tfrac{1}{1}\), \(\tfrac{42}{42}\), \(\tfrac{\sin\theta}{\sin\theta}\), or whatever else over itself, it does not matter. It is still multiplying by 1. You need to keep tat bit of algebra in mind when doing trig. The algebraic manipulations of trig allow you to change the appearance of an equation without changing the actual equality.

OpenStudy (sleepyjess):

Wow

OpenStudy (sleepyjess):

I think I made a mistake while changing the denominator

OpenStudy (e.mccormick):

Yah, which is what my corrected version addressed. When you did it, you did not multiply by 1.

OpenStudy (sleepyjess):

Yes, I accidentally did the first fraction times 1+sinx/cos x and the second one times cos x/1+sin x

OpenStudy (sleepyjess):

When I redo it, I got \(\sf\dfrac{cos^2x}{(1+sin~x)(cos~x)}\) and \(\sf\dfrac{1+sin^2x}{(cos~x)(1+sin~x)}\)

OpenStudy (e.mccormick):

Yep. Anther thing to keep in mind with this sort of problem is the goal, 2sec x. Well, sec x is 1/cos x. So you know that you need to end up with a cos on the bottom. That means that the 1 + sin x on the bottom must canel out somehow. So leave them factored like that. OK. So, do the addition from that point and what does it become?

OpenStudy (e.mccormick):

And I think you have a mistake on the right fraction. On the top.

OpenStudy (sleepyjess):

\(\sf\dfrac{1+cos^2x+sin^2x+2~sin~x}{2((1+sin~x)(cos~x))}\)

OpenStudy (e.mccormick):

\((a + b)(a + b)=a^2+2ab+b^2\) Apply that rule or foil the top of the right and try again. =)

OpenStudy (e.mccormick):

Yah. Tht fixes most of it... but how did you get a 2 on the bottom?

OpenStudy (sleepyjess):

haha yeah, I just added it in on that one, but it should be \(\sf\dfrac{sin^2x+2~sin~x+1}{(1+sin~x)(cos~x)}\)

OpenStudy (sleepyjess):

Whoops, I forgot you don't add the denominators

OpenStudy (sleepyjess):

\(\sf\dfrac{1+cos^2x+sin^2x+2~sin~x}{(1+sin~x)(cos~x)}\)

OpenStudy (e.mccormick):

Yah. OK. So, now you have one fraction, but more still needs to go. Do you see anything else that can be gotten rid of or replaced via an identity?

OpenStudy (sleepyjess):

\(\sf cos^2x+sin^2x=1\)

OpenStudy (e.mccormick):

=) Yep!

OpenStudy (sleepyjess):

So now it would be \(\sf\dfrac{2+2~sin~x}{(1+sin~x)(cos~x)}\)

OpenStudy (e.mccormick):

Yep! Now...see the last thing you need to get rid of?

OpenStudy (sleepyjess):

1+sin x?

OpenStudy (e.mccormick):

Yep. Can you think of what needs to be done to have that on both the top and bottom so that it cancels?

OpenStudy (e.mccormick):

Welp, I need to head out, but at this point you are just a couple of simple steps away from done. =)

OpenStudy (sleepyjess):

:(

OpenStudy (e.mccormick):

Yah, well, there were some service calls. Did you figure out the last steps? It is more algebra than trig at that point. Trig uses a ton of algebra. Calculus uses a ton of EVERYTHING before it. I mean all the way down to basic addition and subtraction on up to the most complex trig identity swaps. So it is good to keep up on earlier math to make later math easier to do.

OpenStudy (sleepyjess):

No, I wasn't sure how to do the rest :/

OpenStudy (e.mccormick):

Well, you have 2+2sx on the top but need the 1+sx on the bottom to cancel. How can you mathematically make the top look more like the bottom?

OpenStudy (sleepyjess):

\(\sf\dfrac{2+sin~x}{cos~x}\)?

OpenStudy (e.mccormick):

Ummm..... I am not sure what you did there. Think of it as an algebra problem at this point. \(\dfrac{2+2a}{(1+a)b}\) It is really the same thing as that. Just did a little substitution. How can I cancel the term (1+a)? What do I need to do to the top to make it work?

OpenStudy (sleepyjess):

\(\sf\dfrac{2(1+sin~x)}{(1+sin~x)(cos~x)}\)

OpenStudy (e.mccormick):

=) OK. So, now it can cancle. That gets you very close to done.

OpenStudy (sleepyjess):

Ok, so after it cancels it is just \(\sf\dfrac{2}{cos~x}\) right?

OpenStudy (e.mccormick):

Yep! Now, do you see how close that is to \(2\sec x\)?

OpenStudy (sleepyjess):

That would be equal to \(\sf 2sec~x\)

OpenStudy (e.mccormick):

Yah. Cause the 2 pops off the top, leaving \(2\dfrac{1}{\cos x}\)

OpenStudy (sleepyjess):

:)

OpenStudy (sleepyjess):

Yah, well, there were some service calls. service calls?

OpenStudy (e.mccormick):

I know this one was a bit of a process. You migh want to see if you can make some clean notes on how it worked. That way it will be more likey to make sense if you see something similar at another time.

OpenStudy (sleepyjess):

Thanks :D

OpenStudy (e.mccormick):

And yes, service calls. As in get in a truck, drive places, and see why their internet is not working.

OpenStudy (sleepyjess):

hmmmm sounds interestina

OpenStudy (sleepyjess):

*interesting

OpenStudy (e.mccormick):

Yah, well, you should see the view if it is on a tower.